# Video-Enhanced Offline Reinforcement Learning: A Model-Based Approach

### v1

@@ Abstract
Offline reinforcement learning (RL) enables policy optimization in static datasets, avoiding the risks and costs of real-world exploration. However, it struggles with suboptimal behavior learning and inaccurate value estimation due to the lack of environmental interaction. In this paper, we present Video-Enhanced Offline RL (VeoRL), a model-based approach that constructs an interactive world model from diverse, unlabeled video data readily available online. Leveraging model-based behavior guidance, VeoRL transfers commonsense knowledge of control policy and physical dynamics from natural videos to the RL agent within the target domain. Our method achieves substantial performance gains (exceeding $100\%$ in some cases) across visuomotor control tasks in robotic manipulation, autonomous driving, and open-world video games. Project page: https://panmt.github.io/VeoRL.github.io.
@@ 1 Introduction
Figure 1: Overview of the training setup of VeoRL that leads to improved offline RL performance. a, We extract latent behavior abstractions from task-agnostic, unannotated natural video data to enrich the world model’s commonsense understanding of the physical world. By interacting with this world model, the agent performs policy optimization guided explicitly by the latent policies learned from the natural videos. b, Overall performance of VeoRL. It demonstrates significant improvements over existing offline RL methods across different visuomotor control benchmarks, including Meta-World robotic manipulation, CARLA autonomous driving, and Minecraft open-world video games. We present the average performance on all tasks on each benchmark.
The core insight of VeoRL is to leverage unlabeled Internet videos to broaden the agent’s understanding of human behaviors and the real physical world, going beyond the limited scope of fixed, predefined datasets. Notably, the vast volume of online video data possesses several key advantages: (i) It can be acquired at a significantly lower cost compared to data collected through real interactions with the target environments. (ii) It implicitly contains demonstration policies that, while potentially suboptimal, can still be beneficial for current visuomotor control tasks.
@@
We evaluate VeoRL on a diverse set of visuomotor control benchmarks. For the Meta-World robotic manipulation tasks, we use the BridgeData-V2 dataset (Walke et al., 2023) as the source of auxiliary real-world videos. For the CARLA autonomous driving tasks, we employ the NuScenes dataset (Caesar et al., 2019), a collection of $1{,}000$ diverse real-world driving scenes, as the auxiliary source domain. For the MineDojo open-world gaming environment, where the agent must navigate a vast state space, we use online Minecraft videos created by human players to provide unlabeled video demonstrations. As summarized in Figure 1b, VeoRL achieves substantial performance improvements over existing RL approaches.
@@ 2 Related Work
For visuomotor control tasks, agents learn control policies from high-dimensional visual inputs. By explicitly modeling state transitions and predicting behavioral outcomes, model-based RL approaches (Ha & Schmidhuber, 2018; Hafner et al., 2019, 2020) typically offer better sample efficiency than model-free methods (Yarats et al., 2021b; Kostrikov et al., 2021; Laskin et al., 2020). World Models (Ha & Schmidhuber, 2018) were introduced to first learn compressed latent representations of the environment in a self-supervised manner, and then use these latent states to train the agent. Dreamer (Hafner et al., 2020) and DreamerV2 (Hafner et al., 2021) optimize behavior by predicting and then acting upon expected values over the latent states generated by the Recurrent State-Space Model. Recently, research on goal-conditioned RL using hierarchical world model architectures (Veeriah et al., 2021; Rao et al., 2022; Hafner et al., 2022; Gumbsch et al., 2023; Park et al., 2024) has shown how decomposing tasks into subgoals can facilitate efficient decision-making in complex environments. However, none of these methods address the challenges of offline RL or facilitate unsupervised policy transfer from enriched out-of-domain video data.
@@ 3.1 Latent Behavior Abstraction
To address the challenge of video data lacking action labels, previous methods (Baker et al., 2022; Zhang et al., 2022) typically focus on using target domain observations with real action labels (i.e., the offline RL dataset) to train an inverse dynamics model, which is then used to assign pseudo-action labels to the source domain auxiliary videos. However, this approach assumes that the source and target domains share the same action space, leading to two potential issues: First, action predictions for source video data may suffer from domain distribution discrepancies. Second, the learned actions are limited in their ability to provide high-level behavior guidance for subsequent policy learning. Instead, we propose a fully unsupervised approach for behavior inference, leveraging a compact categorical latent action space significantly smaller than the raw continuous action space. This approach avoids the challenge of real action prediction, generating higher-level behavior representations that are distinct from the original actions for low-level visuomotor control.
@@ 4 Experiments
In this section, we present (i) quantitative comparisons with existing RL methods on a diverse set of visuomotor control benchmarks, (ii) offline-to-online transfer learning results on novel interactive control tasks, (iii) ablation studies for each proposed model component in VeoRL, and (iv) hyperparameter analyses along with visualizations of the learned latent behavior abstractions. The hyperparameters for implementing VeoRL are provided in Appendix Table 1.
@@ 4.1 Main Results
Figure 6: Performance of offline-to-online finetuning on novel tasks. By conducting offline pretraining, VeoRL demonstrates improved training efficiency and final performance when adapting to unseen, more challenging visuomotor control tasks. Notably, we train the DreamerV2 model under the same offline-to-online setup. The results are averaged over $3$ random training seeds.
@@
To evaluate the generalizability of our model, we finetune the pretrained offline RL agents in unseen visuomotor control tasks through online interactions with the environment. We select a more challenging Soccer task as our offline-to-online transfer target on the Meta-World benchmark. This task typically involves a two-stage decision-making process where the agent must first fetch a ball and then push it to a goal position. For CARLA, the offline RL agents are pretrained using the Day-mode dataset and then finetuned in the Night mode driving scenarios. For MineDojo, the agents are pretrained on a fixed dataset for the Harvest sand task and finetuned on the Harvest cobblestone with wooden pickaxe task. As shown in Figure 6, VeoRL demonstrates superior training efficiency compared to DreamerV2, which follows the same offline-to-online domain adaptation setup, highlighting the effectiveness of our offline training approach.
Figure 7: Ablation studies and correspondence between latent behaviors and real actions. a, We show the impact of latent behavior guidance (yellow) and the intrinsic reward (green). We also compare VeoRL with a baseline model that does not use any source video data throughout the training process (blue). b, During model-based state rollouts, starting from the same state $s_{t}$ of the target domain, VeoRL simultaneously generates real action trajectories $a_{t:t+H}$ and latent behaviors $\bar{a}_{t:t+H}$. We consider trajectories of four consecutive real actions $a_{t:t+4}$, where the latent behaviors are time-invariant, i.e., $\bar{a}_{t}=\bar{a}_{t:t+4}$. We then use t-SNE to project $a_{t:t+4}$ into a 2D plane. To distinguish between different latent behaviors, we assign distinct colors to the points, visualizing $10$ latent behaviors for clarity. We present BridgeData examples where VeoRL generates the corresponding latent behaviors.
@@ 4.2 Ablation Studies
As we will discuss in Section 3, VeoRL introduces three key innovations to the model-based actor-critic learning framework of DreamerV2 (Hafner et al., 2021): (i) It uses latent behaviors as inputs to the policy network (i.e., actor); (ii) It employs latent behaviors to compute an intrinsic reward for training the value network (i.e., critic), and (iii) It trains the plan net in the world model using the auxiliary video set and the offline RL dataset jointly. Accordingly, we conduct ablation studies to evaluate the effectiveness of individual techniques. Figure 7a provides corresponding results on the Handle Press task from Meta-World. Additional results on the other two benchmarks are provided in Appendix Figure 8. As illustrated by the yellow and green bars, excluding the latent behavior input in the policy network or removing its influence from the value network’s training objective both result in significant performance drops in VeoRL. These results highlight the crucial role of latent behaviors as effective high-level guidance in policy learning. The blue bar illustrates that training all network components solely on the target offline dataset (without access to the auxiliary video set) leads to a significant performance decline. This highlights that our video-enhanced method effectively extracts valuable prior knowledge of underlying expert behaviors from the unlabeled videos.
@@ 4.3 Analyses of Latent Actions
To illustrate the correspondence between latent behaviors and real actions, we evaluate the distribution of the learned latent behavior abstractions in Meta-World Coffee Push in Figure 7b. In VeoRL, we roll out the trunk net and plan net simultaneously for model-based policy learning. Starting from the same input state $s_{t}$, we obtain two future trajectories of real actions $a_{t:t+H}$ and corresponding latent behaviors $\bar{a}_{t:t+H}$. To show the correspondence between them, we specifically select trajectories of four consecutive real actions $a_{t:t+4}$ with time-invariant latent behaviors $\bar{a}_{t}=\bar{a}_{t:t+4}$, and then use the t-SNE algorithm (Van der Maaten & Hinton, 2008) to project $a_{t:t+4}$ into a 2D plane. Different colors assigned to the points indicate distinct latent behaviors. From Figure 7b, we observe that the latent behaviors align well with distinct clusters of real action sequences, demonstrating the high representational fidelity of the learned latent action space. These latent behaviors effectively capture semantic information from online videos and can be interpreted as fundamental operational skills, which can then be used as policy guidance for the RL agent. Furthermore, we provide the ablation results of varying the number of discrete latent behavior in Appendix Figure 9.
@@ 5 Conclusions and Limitations
In this paper, we introduced VeoRL, a novel approach that leverages freely available, unlabeled video data to enhance offline visual RL. VeoRL constructs a world model with two state transition branches: one conditioned on latent behaviors extracted from video data and the other on the agent’s real actions. This dual-branch structure allows the agent to effectively learn control policies by aligning state rollouts from two branches, thereby transferring knowledge from the diverse, out-of-domain video data to the RL agents. Our results demonstrate that VeoRL significantly outperforms existing offline RL methods across a wide range of challenging visuomotor control tasks, including robotic manipulation, autonomous driving, and open-world video games. The integration of video data not only boosts the agent’s ability to generalize across different domains but also accelerates the learning process by incorporating a broader understanding of the physical world and control policies.
@@ Appendix C Benchmark Details
Figure 8: Ablation studies in CARLA and MineDojo. We show the impact of latent behavior guidance (yellow) and the intrinsic reward (green). We also compare VeoRL with a baseline model that does not use any source video data throughout the training process (blue).
Appendix D Compared Methods
Appendix E Further Results of Performance Comparison and Ablation Studies
@@
Figure 8 presents the complete ablation results for CARLA and MineDojo.
Appendix F Hyperparameter Analysis of Latent Action Space
Methods DreamerV2 DreamerV3 VeoRL(DV2) VeoRL(DV3)
Success Rate
Drawer Open 0.18 $\pm$ 0.04 0.00 $\pm$ 0.00 0.70 $\pm$ 0.07 0.55 $\pm$ 0.15
Handle Press 0.33 $\pm$ 0.11 0.05 $\pm$ 0.05 0.60 $\pm$ 0.12 0.35 $\pm$ 0.15
Episode Return
Appendix G Comparison with DreamerV3
The decision to use DreamerV2 as the backbone stems from two key considerations: 1) Offline setting performance. While DreamerV3 achieves strong performance across diverse tasks with fixed hyperparameters, our experiments (as shown Table 3) demonstrate that it underperforms in offline RL settings without hyperparameter tuning compared to DreamerV2. 2) Backbone consistency. LAMPO, a key baseline in our work, employs DreamerV2 as its backbone. To enable direct and equitable comparisons between our method and LAMPO, we maintained consistency by using the same backbone.
Notably, we have conducted additional experiments on Meta-world by integrating VeoRL with DreamerV3. The results show that our approach outperforms vanilla DreamerV3, showing its ability to seamlessly integrate with different model architectures.
Appendix H Analyses of the quality of source videos and target domain trajectories
@@
Appendix I Generalizability across tasks
We conduct the experiments on Meta-World and MineDojo to validate the transferability of latent behaviors across tasks. In training phase, the behavior abstraction network is frozen after training on Task A, with a fixed latent codebook. For Task B, this pre-trained network is directly deployed without undergoing task-specific fine-tuning on source video data. As presented in Table 6, the minimal performance degradation confirms that our method’s latent behaviors trained on Task A remain effective for Task B, consistently outperforming other baselines trained specifically on Task B. This transferability significantly reduces training costs for downstream tasks, as it spares the time for re-training on the auxiliary videos.
@@ Appendix I Generalizability across tasks
Appendix J Analyses of MMD loss in domain adaptation
We conducted an ablation study on the MMD loss in the Meta-World environment. To quantify its impact, we compared performance with and without MMD loss, and performed feature distribution visualization using PCA. The results in Table 7 validate that MMD loss critically bridges domain discrepancies. By aligning feature distributions, it enables the target domain to effectively transfer and adapt knowledge from the source domain, leading to a $33\%$ improvement in success rates ($0.60$ vs. $0.45$). Furthermore, as shown in the Figure 10, incorporating MMD significantly aligns the feature distributions between the source and target domains. This alignment correlates with policy generalization.
Table 7: The ablation study on the MMD loss in the Meta-World environment. It shows that removing the MMD loss reduces the performance of our model.

### v2

@@ Abstract
Offline reinforcement learning (RL) enables policy optimization using static datasets, avoiding the risks and costs of extensive real-world exploration. However, it struggles with suboptimal offline behaviors and inaccurate value estimation due to the lack of environmental interaction. We present Video-Enhanced Offline RL (VeoRL), a model-based method that constructs an interactive world model from diverse, unlabeled video data readily available online. Leveraging model-based behavior guidance, our approach transfers commonsense knowledge of control policy and physical dynamics from natural videos to the RL agent within the target domain. VeoRL achieves substantial performance gains (over $100\%$ in some cases) across visual control tasks in robotic manipulation, autonomous driving, and open-world video games. Project page: https://panmt.github.io/VeoRL.github.io.
@@ 1 Introduction
Figure 1: Overview of the training setup of VeoRL that leads to improved offline RL performance. a, We extract latent behavior abstractions from task-agnostic, unannotated natural video data to enrich the world model’s commonsense understanding of the physical world. By interacting with this world model, the agent performs policy optimization guided explicitly by the latent policies learned from the natural videos. b, Overall performance of VeoRL. It demonstrates significant improvements over existing offline RL methods across different visual control benchmarks, including Meta-World robotic manipulation, CARLA autonomous driving, and Minecraft open-world video games. We present the average performance on all tasks on each benchmark.
The core insight of VeoRL is to leverage unlabeled Internet videos to broaden the agent’s understanding of human behaviors and the real physical world, going beyond the limited scope of fixed, predefined datasets. Notably, the vast volume of online video data possesses several key advantages: (i) It can be acquired at a significantly lower cost compared to data collected through real interactions with the target environments. (ii) It implicitly contains demonstration policies that, while potentially suboptimal, can still be beneficial for current visual control tasks.
@@
We evaluate VeoRL on a diverse set of visual control benchmarks. For the Meta-World robotic manipulation tasks, we use the BridgeData-V2 dataset (Walke et al., 2023) as the source of auxiliary real-world videos. For the CARLA autonomous driving tasks, we employ the NuScenes dataset (Caesar et al., 2019), a collection of $1{,}000$ diverse real-world driving scenes, as the auxiliary source domain. For the MineDojo open-world gaming environment, where the agent must navigate a vast state space, we use online Minecraft videos created by human players to provide unlabeled video demonstrations. As summarized in Figure 1b, VeoRL achieves substantial performance improvements over existing RL approaches.
@@ 2 Related Work
For visual control tasks, agents learn control policies from high-dimensional visual inputs. By explicitly modeling state transitions and predicting behavioral outcomes, model-based RL approaches (Ha & Schmidhuber, 2018; Hafner et al., 2019, 2020) typically offer better sample efficiency than model-free methods (Yarats et al., 2021b; Kostrikov et al., 2021; Laskin et al., 2020). World Models (Ha & Schmidhuber, 2018) were introduced to first learn compressed latent representations of the environment in a self-supervised manner, and then use these latent states to train the agent. Dreamer (Hafner et al., 2020) and DreamerV2 (Hafner et al., 2021) optimize behavior by predicting and then acting upon expected values over the latent states generated by the Recurrent State-Space Model. Recently, research on goal-conditioned RL using hierarchical world model architectures (Veeriah et al., 2021; Rao et al., 2022; Hafner et al., 2022; Gumbsch et al., 2023; Park et al., 2024) has shown how decomposing tasks into subgoals can facilitate efficient decision-making in complex environments. However, none of these methods address the challenges of offline RL or facilitate unsupervised policy transfer from enriched out-of-domain video data.
@@ 3.1 Latent Behavior Abstraction
To address the challenge of video data lacking action labels, previous methods (Baker et al., 2022; Zhang et al., 2022) typically focus on using target domain observations with real action labels (i.e., the offline RL dataset) to train an inverse dynamics model, which is then used to assign pseudo-action labels to the source domain auxiliary videos. However, this approach assumes that the source and target domains share the same action space, leading to two potential issues: First, action predictions for source video data may suffer from domain distribution discrepancies. Second, the learned actions are limited in their ability to provide high-level behavior guidance for subsequent policy learning. Instead, we propose a fully unsupervised approach for behavior inference, leveraging a compact categorical latent action space significantly smaller than the raw continuous action space. This approach avoids the challenge of real action prediction, generating higher-level behavior representations that are distinct from the original actions for low-level visual control.
@@ 4 Experiments
In this section, we present (i) quantitative comparisons with existing RL methods on a diverse set of visual control benchmarks, (ii) offline-to-online transfer learning results on novel interactive control tasks, (iii) ablation studies for each proposed model component in VeoRL, and (iv) hyperparameter analyses along with visualizations of the learned latent behavior abstractions. The hyperparameters for implementing VeoRL are provided in Appendix Table 2.
@@ 4.1 Main Results
Figure 6: Performance of offline-to-online finetuning on novel tasks. By conducting offline pretraining, VeoRL demonstrates improved training efficiency and final performance when adapting to unseen, more challenging visual control tasks. Notably, we train the DreamerV2 model under the same offline-to-online setup. The results are averaged over $3$ random training seeds.
@@
To evaluate the generalizability of our model, we finetune the pretrained offline RL agents in unseen visual control tasks through online interactions with the environment. We select a more challenging Soccer task as our offline-to-online transfer target on the Meta-World benchmark. This task typically involves a two-stage decision-making process where the agent must first fetch a ball and then push it to a goal position. For CARLA, the offline RL agents are pretrained using the Day-mode dataset and then finetuned in the Night mode driving scenarios. For MineDojo, the agents are pretrained on a fixed dataset for the Harvest sand task and finetuned on the Harvest cobblestone with wooden pickaxe task. As shown in Figure 6, VeoRL demonstrates superior training efficiency compared to DreamerV2, which follows the same offline-to-online domain adaptation setup, highlighting the effectiveness of our offline training approach.
Figure 7: Ablation studies of VeoRL. We show the impact of latent behavior guidance (yellow) and the intrinsic reward (green). We also compare VeoRL with a baseline model that does not use any source video data throughout the training process (blue).
Methods DreamerV2 DreamerV3 VeoRL(DV2) VeoRL(DV3)
Success Rate
Drawer Open 0.18 $\pm$ 0.04 0.00 $\pm$ 0.00 0.70 $\pm$ 0.07 0.55 $\pm$ 0.15
Handle Press 0.33 $\pm$ 0.11 0.05 $\pm$ 0.05 0.60 $\pm$ 0.12 0.35 $\pm$ 0.15
Episode Return
@@ 4.2 Ablation Studies
As we will discuss in Section 3, VeoRL introduces three key innovations to the model-based actor-critic learning framework of DreamerV2 (Hafner et al., 2021): (i) It uses latent behaviors as inputs to the policy network (i.e., actor); (ii) It employs latent behaviors to compute an intrinsic reward for training the value network (i.e., critic), and (iii) It trains the plan net in the world model using the auxiliary video set and the offline RL dataset jointly. Accordingly, we conduct ablation studies to evaluate the effectiveness of individual techniques. Figure 7 provides corresponding results on Meta-World (Handle Press), CARLA and MineDojo (Harvest water with bucket). As illustrated by the yellow and green bars, excluding the latent behavior input in the policy network or removing its influence from the value network’s training objective both result in significant performance drops in VeoRL. These results highlight the crucial role of latent behaviors as effective high-level guidance in policy learning. The blue bar illustrates that training all network components solely on the target offline dataset (without access to auxiliary videos) leads to a significant performance decline. This highlights that our video-enhanced method effectively extracts valuable prior knowledge of underlying expert behaviors from the unlabeled videos.
4.3 Comparison with DreamerV3
The decision to use DreamerV2 as the backbone is driven by two key considerations: (i) Offline setting performance. While DreamerV3 (Hafner et al., 2023) achieves strong performance across diverse tasks with fixed hyperparameters, our experiment results in Table 1 demonstrate that it underperforms in offline RL settings without hyperparameter tuning compared to DreamerV2. (ii) Backbone consistency. LAMPO, a key baseline in our work, uses DreamerV2 as its backbone. To ensure direct and fair comparisons between our method and LAMPO, we maintain consistency by adopting the same backbone. Notably, we also conduct additional experiments on Meta-World by integrating VeoRL with DreamerV3. The results demonstrate that our approach outperforms vanilla DreamerV3, showcasing its ability to seamlessly integrate with different architectures.
Figure 8: Correspondence between latent behaviors and real actions in Meta-World. During model-based state rollouts, starting from the same state $s_{t}$ of the target domain, VeoRL simultaneously generates real action trajectories $a_{t:t+H}$ and latent behaviors $\bar{a}_{t:t+H}$. We consider trajectories of four consecutive real actions $a_{t:t+4}$, where the latent behaviors are time-invariant, i.e., $\bar{a}_{t}=\bar{a}_{t:t+4}$. We then use t-SNE to project $a_{t:t+4}$ into a 2D plane. To distinguish between different latent behaviors, we assign distinct colors to the points, visualizing $10$ latent behaviors for clarity. We present BridgeData examples where VeoRL generates the corresponding latent behaviors.
@@ 4.4 Analyses of Latent Actions
To illustrate the correspondence between latent behaviors and real actions, we evaluate the distribution of the learned latent behavior abstractions in Meta-World Coffee Push in Figure 8. In VeoRL, we roll out the trunk net and plan net simultaneously for model-based policy learning. Starting from the same input state $s_{t}$, we obtain two future trajectories of real actions $a_{t:t+H}$ and corresponding latent behaviors $\bar{a}_{t:t+H}$. To show the correspondence between them, we specifically select trajectories of four consecutive real actions $a_{t:t+4}$ with time-invariant latent behaviors $\bar{a}_{t}=\bar{a}_{t:t+4}$, and then use the t-SNE algorithm (Van der Maaten & Hinton, 2008) to project $a_{t:t+4}$ into a 2D plane. Different colors assigned to the points indicate distinct latent behaviors. From Figure 8, we observe that the latent behaviors align well with distinct clusters of real action sequences, demonstrating the high representational fidelity of the learned latent action space. These latent behaviors effectively capture semantic information from online videos and can be interpreted as fundamental operational skills, which can then be used as policy guidance for the RL agent. Furthermore, we provide the ablation results of varying the number of discrete latent behavior in Appendix Figure 9.
4.5 Further Results and Discussion
Please refer to Appendices F-I for additional experimental results, including the quality analysis of source videos and target domain trajectories, the hyperparameter analysis of latent action space, the generalizability across tasks, and the impact of MMD in domain adaptation.
@@ 5 Conclusions and Limitations
In this paper, we introduced VeoRL, a novel approach that leverages freely available, unlabeled video data to enhance offline visual RL. VeoRL constructs a world model with two state transition branches: one conditioned on latent behaviors extracted from video data and the other on the agent’s real actions. This dual-branch structure allows the agent to effectively learn control policies by aligning state rollouts from two branches, thereby transferring knowledge from the diverse, out-of-domain video data to the RL agents. Our results demonstrate that VeoRL significantly outperforms existing offline RL methods across a wide range of challenging visual control tasks, including robotic manipulation, autonomous driving, and open-world video games. The integration of video data not only boosts the agent’s ability to generalize across different domains but also accelerates the learning process by incorporating a broader understanding of the physical world and control policies.
@@ Appendix D Compared Methods
Appendix E Further Results of Performance Comparison
@@
Appendix F Analyses of the Quality of Source Videos and Target Domain Trajectories
@@
Appendix G Hyperparameter Analysis of Latent Action Space
Appendix H Generalizability across Tasks
We conduct the experiments on Meta-World and MineDojo to validate the transferability of latent behaviors across tasks. In the training phase, the behavior abstraction network is frozen after training on Task A, with a fixed latent codebook. For Task B, this pre-trained network is directly deployed without undergoing task-specific fine-tuning on source video data. As presented in Table 6, the minimal performance degradation confirms that our method’s latent behaviors trained on Task A remain effective for Task B, consistently outperforming other baselines trained specifically on Task B. This transferability significantly reduces training costs for downstream tasks, as it spares the time for re-training on the auxiliary videos.
@@ Appendix H Generalizability across Tasks
Appendix I Analyses of MMD in Domain Adaptation
We performed an ablation study in the Meta-World environment to evaluate the role of MMD loss. To quantify its impact, we compared performance under conditions with and without MMD loss, accompanied by feature distribution visualizations derived from principal component analysis (PCA). The results in Table 7 validate that MMD loss critically bridges domain discrepancies. By aligning feature distributions, it enables the target domain to effectively transfer and adapt knowledge from the source domain, leading to a $33\%$ improvement in success rates ($0.60$ vs. $0.45$). Furthermore, as shown in the Figure 10, incorporating MMD significantly aligns the feature distributions between the source and target domains. This alignment correlates with policy generalization.
Table 7: The ablation study on the MMD loss in the Meta-World environment.